# Dynamic Routing Between Experts: A Data-Efficient Approach to Continual Learning in Vision-Language Models

### v2

@@ Abstract
Vision-Language Models (VLMs) suffer from catastrophic forgetting when sequentially fine-tuned on new tasks, degrading performance on previously learned foundational and task-specific capabilities. While multi-task learning can mitigate forgetting, it requires simultaneous access to all datasets and imposes computational overhead that scales linearly with the number of tasks. In this work, we introduce a routing-based approach that enables the integration of new tasks while preserving the foundational knowledge acquired during pretraining. We evaluate our method using InternVL-2 models (2B and 8B parameters) and demonstrate that routing preserves the model’s foundational capabilities by maintaining performance on general-purpose benchmarks such as ChartQA, MMBench, and DocVQA, while simultaneously improving accuracy on specialized tasks. Importantly, our approach achieves this without requiring concurrent access to data from all tasks, avoiding the significant computational and data overhead associated with traditional multi-task learning. We further conduct extensive ablation studies to evaluate the scalability and robustness of routing-based learning, showing that the approach is resilient to a growing number of tasks and performs particularly well when new tasks are semantically related. Finally, we show that the routing mechanism enables superior cross-modal transfer between language and vision capabilities, allowing knowledge learned in one modality to enhance performance in another capability not achieved by existing continual learning methods.
Introduction
Figure 1: Comparison of Multi-Task Learning, Routing-Based Methods and Sequential FT in mitigating catastrophic forgetting. Routing-based methods match multi-task learning performance while requiring only new task data, avoiding costly retraining on previous tasks. Tasks are introduced sequentially from left to right, with the multi-task model trained on the cumulative data from all prior tasks. In contrast, sequential fine-tuning only sees new task data, leading to strong performance on the most recent task but potential forgetting of earlier ones.
Large Language Models (LLMs) have advanced the state-of-the-art in Natural Language Processing (NLP), demonstrating capabilities in tasks such as machine translation, sentiment analysis, and question answering (Brown et al. 2020; Chowdhery et al. 2023). Building on advances in language modeling and computer vision, Vision-Language Models (VLMs) have emerged as multimodal systems capable of joint reasoning over visual and textual inputs (Radford et al. 2021; Lu et al. 2019), enabling applications from image captioning to visual question answering (Chen et al. 2020).
Although pretrained VLMs offer strong general-purpose capabilities, fine-tuning them on specialized-domain tasks (such as medical imagery, legal documents) is often required to achieve top performance on those specialized domains. However, this fine-tuning frequently triggers Catastrophic Forgetting (CF)—a phenomenon where the model starts to loose general purpose capabilities acquired from large-scale multimodal corpora (Kirkpatrick et al. 2017; Kemker et al. 2018; Goodfellow et al. 2013).
Recent work shows that CF severity increases with model scale (Luo et al. 2023), making this particularly challenging for LLMs. Kalajdzievski (2024) et al. further establishes precise scaling laws: the extent of forgetting scales predictably with both the number of fine-tuned parameters and the number of update steps, even under parameter-efficient paradigms such as Low-Rank Adaptation (LoRA) (Hu et al. 2022). In multimodal settings, (Zhai et al. 2023) demonstrate that fine-tuning large VLMs on one modality (e.g., visual tasks) degrades performance in the other (e.g., language tasks), reinforcing the interference between modalities during sequential adaptation.
A common strategy to mitigate CF is multi-task learning, where all tasks are trained jointly to preserve shared representations (Ruder 2017). While this setup often serves as an upper bound for continual learning performance—since it assumes access to all task data—it is computationally expensive and infeasible in many real-world settings due to the need for concurrent dataset access and large-scale infrastructure.
Continual learning offers a practical alternative to multi-task learning and has been explored through several strategies, including replay-based methods (Rolnick et al. 2019; Shin et al. 2017), regularization techniques (Kirkpatrick et al. 2017; Zenke, Poole, and Ganguli 2017), knowledge distillation (Li and Hoiem 2017; Rebuffi et al. 2017), and parameter-isolation approaches (Rusu et al. 2016; Mallya and Lazebnik 2018; Wortsman et al. 2020; Mallya, Davis, and Lazebnik 2018; Mallya and Lazebnik 2022). While effective in smaller-scale models or unimodal settings, these methods often require access to previous task data or model checkpoints, or introduce architectural changes that hinder scalability to billion-parameter vision-language models (VLMs) (Chen et al. 2024b; Liu et al. 2023). To address these scalability challenges, recent work SEMA (Wang et al. 2024a) has turned to adapter-based techniques, which offer a modular and lightweight alternative by enabling task-specific tuning without altering the core model. While SEMA supports sequential training without requiring access to all task data simultaneously, it remains limited to single-modality settings and coarse layer-level adaptation.
To address the scalability and data access limitations of traditional continual learning methods, model merging has emerged as a promising modular alternative. These techniques combine the weights of separately fine-tuned models to integrate new capabilities into a base model—without requiring concurrent access to all task data (Matena and Raffel 2022; Douillard et al. 2021; Ilharco et al. 2023; Merlingot et al. 2024; Cha et al. 2020). By operating directly on model parameters, merging offers an efficient way to incorporate task-specific knowledge while avoiding catastrophic forgetting. This makes it particularly attractive for large-scale vision-language models, where full retraining or replay is infeasible. However, merging often underperforms compared to joint multi-task training, especially when tasks are semantically distant or involve modality-specific reasoning, leading to interference and degraded performance in the merged model.
Another family of modular approaches—routing-based methods focuses on dynamically selecting specialized sub-networks based on input characteristics (Li et al. 2018; Ostapenko et al. 2024; Muqeeth et al. 2024). While conceptually related to Mixture of Experts (MoE) architectures, these methods differ in both design goals and constraints. MoE models for vision-language tasks, such as MoE-LLaVA (Lin et al. 2024) and V-MoE (Riquelme et al. 2021), primarily aim to enhance model capacity and computational efficiency through parallel expert networks activated by learned gating mechanisms (Shazeer et al. 2017; Fedus, Zoph, and Shazeer 2021). However, MoE architectures typically require simultaneous access to all datasets during training and focus on improving zero-shot generalization rather than preserving previously acquired knowledge. Similarly, existing routing-based approaches have predominantly been developed to improve generalization to novel inputs, leaving a significant gap in their application to continual learning scenarios where maintaining performance on previous tasks is essential.
In this work, we adapt routing-based modularity specifically to address catastrophic forgetting in large-scale vision-language models. We systematically investigate token-level routing within large-scale vision-language models as a strategy to overcome forgetting. Unlike traditional multi-task learning, which relies on joint training over all datasets, our approach supports incremental learning using only the current task’s data—while achieving performance comparable to multi-task baselines, illustrated in Figure 1. We evaluate the approach across multiple sequential tasks and analyze its scalability with increasing model size. Our experiments demonstrate that as model size grows, the performance gap between our routing method and fully specialized models narrows. Furthermore, we highlight the unique ability of token-level routing to facilitate cross-modal transfer, enabling knowledge gained in one modality to enhance performance in another—a critical capability not achieved by existing continual learning methods.
Preliminaries
We first review foundational concepts relevant to continual learning in vision-language models, including the underlying model architecture and common continual learning strategies such as multi-task learning, sequential fine-tuning and expert merging.
Vision Language Models
Vision-Language Models (VLMs) process multimodal inputs by encoding visual content through a visual encoder $V$ and projecting these features into a shared embedding space via projector $F$, where visual features and textual embeddings are fused and jointly processed by a transformer-based language model for multimodal reasoning.
Multi-task Learning
Multi-task learning (MTL) mitigates catastrophic forgetting by training on all specialized tasks simultaneously, enabling shared representations that generalize across tasks (Caruana 1997). While MTL serves as a performance upper bound for CL, it is often computationally prohibitive due to simultaneous access and training on all task data.
Sequential Fine-tuning
Sequential fine-tuning adapts pre-trained models to new tasks incrementally, making it computationally efficient but highly susceptible to catastrophic forgetting (Goodfellow et al. 2013). Techniques like experience replay (Rolnick et al. 2019) improve knowledge retention but require storing previous data, limiting their practicality for large-scale applications. Other approaches to mitigate catastrophic forgetting include regularization-based methods (Kirkpatrick et al. 2017; Zenke, Poole, and Ganguli 2017) and parameter isolation techniques (Rusu et al. 2016; Mallya and Lazebnik 2018), but these often involve architectural modifications or require access to prior task data, making them challenging to scale to billion-parameter vision-language models.
Expert Merging
Model merging combines multiple task-specific models into a single aggregate model. Starting with common architecture and initial weights $\theta_{0}$, merging integrates specialized knowledge from models with weights $\theta_{1},\theta_{2},\ldots,\theta_{n}$. We explore two strategies:
Model averaging (McMahan et al. 2017) uses an element-wise average of model parameters: $\theta_{m}=\frac{1}{n}\sum_{i=0}^{n}\theta_{i}$
Task Arithmetic (Ilharco et al. 2023) constructs task vectors and combines them with the base model: $\theta_{m}=\theta_{0}+\sum_{i=1}^{n}\lambda_{i}t_{i}$ where $t_{i}=\theta_{i}-\theta_{0}$
These methods require only the dataset of the new specialized task, making them cost-effective alternatives for mitigating catastrophic forgetting.
Approach
Routing between experts is a modular machine learning paradigm that leverages a routing function to dynamically direct inputs to specialized expert modules based on input embeddings. Prior works such as PHATGOOSE (Muqeeth et al. 2024) and Arrow (Ostapenko et al. 2024) have primarily explored routing to enable zero-shot generalization in large language models by dynamically selecting expert modules at inference.
While most routing algorithms have been explored in the context of LLMs, we adapt them for VLMs with a different objective: mitigating CF rather than enabling zero-shot generalization. Building on the superior performance of PHATGOOSE (Muqeeth et al. 2024), we extend its design by incorporating specialized LoRA modules tailored to VLM architectures, and implement a routing function that dynamically selects task-specific adapters during inference. Below, we detail how PHATGOOSE’s routing mechanism is applied within the vision-language domain.
PHATGOOSE for VLMs. Building on the VLM architecture described previously, we implement task-specific LoRA adapters within the language model component while keeping the vision encoder $V$ and projection layer $F$ frozen. This design choice enables a more controlled evaluation of catastrophic forgetting, isolating the impact of adaptation within the language model and minimizing confounding factors.
@@ Introduction
Setup. In this setup, LoRA modifies the output of each linear layer in the language model while keeping the base parameters frozen. Specifically, for a linear layer with weights $W\in\mathbb{R}^{d\times n}$, given an input activation $u_{t}\in\mathbb{R}^{n}$, the output is computed as:
$Wu_{t}+BAu_{t},$
where $A\in\mathbb{R}^{r\times n}$ and $B\in\mathbb{R}^{d\times r}$ are the trainable LoRA parameters, while $W$ remains frozen during fine-tuning. Here, $u_{t}$ denotes the input activation corresponding to the $t$-th token in the sequence. The dimensions $d$ and $n$ represent the output and input sizes of the linear layer, respectively, and $r\ll\min(d,n)$ is a small rank hyperparameter that controls the number of trainable parameters.
Training. Our training procedure consists of two sequential stages designed to enable modular adaptation without modifying the base model or requiring access to previous tasks. First, we train task-specific LoRA modules that specialize the language model for each new task independently. During this stage, all base model parameters, including the vision encoder and core language model weights, remain frozen. Only the LoRA parameters are updated using data exclusively from the current task. This isolates task adaptation and prevents interference.
After the LoRA modules are trained, we introduce lightweight routing vectors—one per task—that serve as gating mechanisms to dynamically control the influence of each LoRA module during inference. Crucially, each routing vector is trained independently using only the task-specific data and the corresponding LoRA module, without accessing other tasks’ data or parameters. Initialized to zero, the routing vector is optimized with the same next-token prediction objective used for LoRA training, while keeping all other parameters—including the base model and all LoRA modules—frozen. Through this process, the routing vector learns to identify input activation patterns unique to its task and applies gating on a token-level basis via a sigmoid function, effectively controlling the task-specific adaptation dynamically.
Specifically, the output of a linear layer augmented with routing and LoRA modules is computed as:
$Wu_{t}+BAu_{t}\sigma(v^{T}u_{t}),$
where $W$ is the frozen base weight matrix, $A$ and $B$ are the LoRA parameters trained per task, $u_{t}$ is the input activation at token $t$, and $\sigma(v^{T}u_{t})$ is the sigmoid gating function controlled by the routing vector $v$.
We assume tasks arrive sequentially. For each new task, only the corresponding LoRA module and routing vector are trained using data from that task alone. During specialization, no data or parameters from previous tasks are accessed or modified, thereby enforcing continual learning constraints and enabling efficient adaptation without catastrophic forgetting.
Inference. To enable dynamic routing during inference, we load all the LoRA modules and gating vectors trained for different tasks. The gating vectors $v_{1},v_{2},\dots,v_{n}$ correspond to tasks $\{1,2,\dots,n\}$. All routing decisions and module activations are computed independently for each token in the input sequence, ensuring fine-grained token-level routing. At each layer, routing proceeds as follows:
@@
  • Compute the affinity score between the input activation $u_{t}$ and each gating vector $v_{i}$ via dot product:
    $\alpha_{t,i}=v_{i}^{T}u_{t}.$
  • Select the top-$k$ modules with the highest affinity scores, where $k$ is a user-defined hyperparameter.
  • Normalize these scores using a softmax scaled by $\sqrt{n}$, where $n$ is the input dimension, to obtain weights:
    $w_{t}=\mathrm{softmax}\left(\left\{\frac{\alpha_{t,i}}{\sqrt{n}}:i\in E_{t}\right\}\right),$
    where $E_{t}$ is the set of selected top-$k$ modules.
The output of the routed linear layer at token position $t$ is then computed as:
$Wu_{t}+\sum_{i\in E_{t}}w_{t,i}B_{i}A_{i}u_{t},$
where $W$ is the frozen base weight matrix, and $A_{i},B_{i}$ are the LoRA parameters for module $i$.
Crucially, this inference design respects continual learning constraints: it does not require access to any previous task data or task identifiers. New tasks can be added simply by training their respective LoRA modules and routing vectors without altering existing modules. The routing vectors implicitly learn to recognize input patterns and gate module activation, thus preventing interference between tasks and mitigating catastrophic forgetting while preserving the base model’s foundational capabilities.
Experiment Design
In this section, we present the models, datasets, training hyperparameters, and strategies employed to evaluate our continual learning framework using routing-based approaches.
Models
We use InternVL-2 (Chen et al. 2024b), one of the most powerful open-source VLM in our experiments. The InternVL-2 family includes models ranging from 1B to 108B parameters. The model supports multimodal input, handling text, images, video. We utilize 2B and 8B versions of InternVL-2 for our experiments on CF.
Datasets and Evaluations
To assess CF across different methods, we utilize a combination of specialization tasks and benchmark tasks. Specialization tasks involve fine-tuning the model to enhance performance on specific tasks, while benchmark tasks evaluate the model’s original proficiency, providing a measure of its general-purpose capability. Achieving strong performance on both specialization and benchmark tasks is critical for any method because that would mean the model gains new knowledge while preserving the older information.
For specializing tasks, we utilize several datasets that cover a diverse range of multimodal applications such as image captioning, visual entailment, hateful content detection, and multimodal classification, offering a comprehensive framework for evaluating the effectiveness and robustness of our models. The COCO-Caption (COCO) dataset (Lin et al. 2015), focused on image captioning, is evaluated using the CIDEr metric. SNLI-VE (SNLI) (Xie et al. 2018), a text-image entailment task, is assessed based on accuracy. The Hateful Memes (HM) dataset (Kiela et al. 2021), which focuses on detecting hateful memes, is also evaluated using accuracy. Finally, we use the MMIMDb dataset (Arevalo et al. 2017) for movie genre classification, with performance measured by precision.
To quantify the model’s retained capabilities post-specialization, we utilize three benchmark tasks. MMBench (MMB) (Liu et al. 2024) evaluates preservation of fine-grained perception and reasoning abilities across 20 dimensions, ChartQA (Masry et al. 2022a) tests retention of data visualization comprehension skills, and DocVQA (Mathew, Karatzas, and Jawahar 2021b) assesses document understanding capabilities. This evaluation strategy enables us to assess both the model’s ability to acquire new specialized knowledge and its capacity to retain previously learned skills.
As our routing-based framework enables zero-shot cross-modal transfer, we conduct experiments training two expert models: one specialized on the text modality using the Orca-Math dataset (Mitra et al. 2024), and another trained on the combined image+text modality using the Multi-math dataset (Peng et al. 2024). For zero-shot evaluation, we employ the MGSM framework (Shi et al. 2022) on both English and Chinese subsets to measure cross-lingual and cross-modal generalization.
Training
The models are trained using a cosine learning rate scheduler to adjust the learning rate over the course of training. For specializing LoRA module, we employ a learning rate of $4\times 10^{-5}$. LoRA modules are added exclusively to the LLM, ensuring that only the LoRA weights of the LLM are updated during training. To train the gating module we freeze all the parameters and only train the gate vector $v$ for 100 steps on the specialized task.
For sequential learning, we follow the task progression: COCO-Caption → SNLI-VE → Hateful Memes → MMIMDb. This sequence allows us to analyze the impact of each approach across diverse datasets and tasks. During inference we set the top-$k$ hyperparameter in $k=1$ for all of our experiments.
Results
In this section, we present the results of our study organized into two main categories: (1) comparative evaluations of different continual learning methods, and (2) ablation studies examining the properties and capabilities of our routing-based approach.
Continual Learning Comparison
InternVL2-2B Results. We evaluate various continual learning strategies on the InternVL2-2B model across multiple tasks. The approaches include MTL training, sequential fine-tuning, experience replay, model merging, and routing.
Method Specialized Tasks Benchmark Tasks
COCO SNLI HM MMIMDb MMB ChartQA DocVQA
Zero-shot 0.795 52.9 60.2 46.6 79.1 59.5 84.7
Sequential FT 1.181 61.0 49.7 62.3 67.1 41.2 70.0
Sequential FT w/ ER (20%) 1.280 58.1 73.8 62.3 70.9 46.6 73.9
Merging: Model Avg. 1.148 72.4 64.1 62.3 78.8 59.2 83.6
Merging: Task Arith. 1.171 73.3 64.5 64.9 78.6 59.2 84.1
Routing 1.305 80.3 71.2 63.2 78.8 59.2 84.3

### v3

# Routing-Based Continual Learning for Multimodal Large Language Models

@@ Abstract
Multimodal Large Language Models (MLLMs) struggle with continual learning, often suffering from catastrophic forgetting when adapting to sequential tasks. We introduce a routing-based architecture that integrates new capabilities while robustly preserving foundational knowledge. While Multi-Task Learning (MTL) offers a theoretical performance upper bound, it incurs a linearly scaling computational overhead as the number of tasks increases. In contrast, our method maintains fixed data and compute requirements regardless of the task sequence length. Across models ranging from 2B to 8B parameters, we demonstrate that our routing approach performs on par with MTL while retaining the training efficiency of sequential fine-tuning. Beyond merely mitigating forgetting, we observe that token-level routing facilitates cross-modal transfer, leveraging knowledge from one modality to bolster performance in another. Ablation studies confirm the approach’s scalability: routing remains robust even with large expert pools and effectively capitalizes on task relatedness. Finally, we show that our method scales favorably, with larger models exhibiting minimal degradation compared to fully specialized fine-tuning.
Routing-Based Continual Learning for Multimodal
Large Language Models
Jay Mohtathanks: Equal contribution , Kenan Emir Ak11footnotemark: 1 , Gwang Lee, Dimitris Dimitriadis, Yan Xu, Mingwei Shen Amazon.com {jaymoht, kenanea, gglee, dbdim, yanxuml, mingweis}@amazon.com
1 Introduction
Large Language Models (LLMs) have fundamentally transformed Natural Language Processing (NLP) Brown et al. (2020), a success now mirrored by Multimodal Large Language Models (MLLMs) capable of reasoning leveraging visual and textual inputs Radford et al. (2021); Alayrac et al. (2022); Liu et al. (2023). However, adapting these general-purpose models to specialized domains introduces a critical dilemma: optimizing performance on target tasks via fine-tuning frequently triggers catastrophic forgetting (CF), severely degrading the model’s previously acquired knowledge McCloskey and Cohen (1989).
This challenge is most severe in continual learning settings, where models must learn a sequence of tasks without access to historical training data Kemker et al. (2018). Given the massive scale of MLLMs, traditional solutions such as experience replay Rolnick et al. (2019); Rebuffi et al. (2017) or parameter regularization Kirkpatrick et al. (2017) are often computationally prohibitive or ineffective Wu et al. (2024). Because general linguistic and visual knowledge is distributed across billions of parameters, task-specific updates may lead to catastrophic forgetting, where the optimization for new objectives inadvertently overwrites the representations necessary for earlier tasks Kotha et al. (2023).
Figure 1: Comparison of continual learning methods. Left: Task performance across four vision-language benchmarks. Right: Training time and inference requirements. Our routing-based approach performs on par with Multi-Task Learning performance (upper bound) while requiring $2.7\times$ less training time and only a single forward pass at inference, unlike MR-LoRA which requires two.
In the multimodal context, this interference is particularly damaging as it risks the cross-modal alignment established during pretraining Zhai et al. (2023). Fine-tuning can decouple visual embeddings from their textual counterparts, breaking the model’s reasoning capabilities Zhu et al. (2024). Furthermore, visual encoders and language decoders often suffer from asymmetric forgetting rates, further destabilizing the model Ni et al. (2023). Consequently, there is a critical need for architectures that allow plasticity for new tasks while enforcing stability for foundational cross-modal alignments.
Figure 1 illustrates the trade-offs between these paradigms. While Sequential Fine-Tuning is data-efficient, it suffers from a sharp decline in performance on prior tasks. Conversely, Multi-Task Learning (MTL) ensures stability by retraining on the cumulative dataset, but incurs a linearly scaling computational cost that is unsustainable for large-scale MLLMs. Furthermore, while modular frameworks like MR-LoRA Zhao et al. (2025) reduce interference, they often require two forward passes at inference to resolve task-specific outputs. Our approach aims to occupy a middle ground: it achieves the stability of MTL and the efficiency of sequential methods while maintaining only a single-pass inference.
In this work, we introduce a routing-based modular architecture designed to tackle the challenges of continual learning in MLLMs. By leveraging token-level routing Muqeeth et al. (2024), our method isolates task-specific adaptations to prevent the overwriting of shared knowledge. Our contributions are fourfold:
  1. 1.
    We demonstrate that routing-based continual learning matches the performance stability of Multi-Task Learning using only the current task’s data, eliminating the need for costly retraining.
  2. 2.
    We uncover that token-level routing enables distinct cross-modal transfer, allowing improvements in unimodal experts (e.g., text-only) to bolster performance in multimodal domains.
  3. 3.
    We provide extensive evaluations across sequential tasks, demonstrating robustness to expert pool size and the positive utilization of task relatedness.
  4. 4.
    We identify a positive scaling trend: as model size increases, our method approaches the performance of fully specialized fine-tuning, validating its viability for large-scale deployment.
2 Related Work
2.1 Multimodal Large Language Models
MLLMs extend the capabilities of LLMs by processing visual inputs alongside text, typically comprising a pretrained vision encoder, a language backbone, and a connector module that projects visual features into the language embedding space Yin et al. (2024). Early architectures, such as Flamingo Alayrac et al. (2022) and BLIP-2 Li et al. (2023), established foundational design principles by bridging frozen vision and language models via modular components.
Subsequent research has prioritized scaling model capacity and refining instruction tuning. LLaVA Liu et al. (2023, 2024a) demonstrated the efficacy of visual instruction tuning using simple projection layers, while models such as Qwen-VL Bai et al. (2023) and InternVL Chen et al. (2024b, a) introduced multi-stage training paradigms incorporating diverse data mixtures and dynamic resolution processing. However, these architectures are typically optimized for static training pipelines under the assumption of i.i.d. data distributions. Consequently, how to effectively incorporate continual adaptation into these frameworks remains unclear; without explicit mechanisms to manage sequential learning, they remain susceptible to catastrophic forgetting.
2.2 Continual Learning in Large Models
Catastrophic forgetting remains a persistent challenge in deep learning, exhibiting a relationship with model scale that is both complex and task-dependent Ramasesh et al. (2021); Kalajdzievski (2024). In MLLMs, this difficulty is intensified by cross-modal interference: sequential adaptation often disrupts the alignment established during pretraining, causing fine-tuning in one modality to degrade representations in the other Zhai et al. (2023).
Traditional continual learning strategies face severe scalability bottlenecks when applied to billion-parameter models. Replay-based methods Rolnick et al. (2019); Chaudhry et al. (2019) require the retention of historical data, raising privacy and storage concerns, while regularization techniques Kirkpatrick et al. (2017); Zenke et al. (2017) frequently demand computationally prohibitive importance scoring across the entire network. Although multi-task learning provides a theoretical upper bound for performance, it necessitates simultaneous access to all datasets, rendering it impractical for dynamic data streams Ruder (2017).
To address these limitations, recent Parameter-Efficient Fine-Tuning (PEFT) methods have been adapted for the continual learning setting. Approaches such as Progressive Prompts Razdaibiedina et al. (2023) utilize task-specific tokens, while O-LoRA Wang et al. (2023) constrains adaptation to orthogonal subspaces. Similarly, SEMA Wang et al. (2025) introduces adapter-based sequential training with automatic expansion for vision-based learning. However, the MLLM-CL benchmark Zhao et al. (2025) indicates that while these approaches mitigate forgetting, they struggle with complex, domain-specific shifts. Unlike these adapter-centric strategies, our method operates at the token level, enabling fine-grained routing that better accommodates vision-language tasks.
2.3 Model Merging and Modular Routing
Modular architectures offer a promising alternative to rigid end-to-end training. Model merging, for instance, combines the weights of separately fine-tuned models without requiring concurrent data access Matena and Raffel (2022); Ilharco et al. (2022); Yadav et al. (2023). While computationally efficient, merging is static; it often underperforms when tasks are semantically distant or require conflicting reasoning patterns, leading to interference in the final averaged weights.
Routing-based methods address these challenges by dynamically selecting specialized sub-networks based on input semantics. While frameworks like PHATGOOSE Muqeeth et al. (2024) and Arrow Ostapenko et al. (2024) utilize modular routing for zero-shot generalization, and Mixture-of-Experts (MoE) architectures Lin et al. (2024); Riquelme et al. (2021) employ it for capacity scaling, their application in continual learning remains underexplored. Specifically, standard MoEs often lack mechanisms to isolate task-specific knowledge or handle the sequential arrival of data. Our approach addresses this by leveraging routing to manage an ever-growing library of Low-Rank Adapters (LoRA). Rather than updating a static set of parameters, we instantiate a new LoRA ’specialist’ for each incremental task. The router then learns to steer inputs toward the appropriate module, allowing the MLLM to broaden its capabilities while maintaining strict functional isolation to prevent catastrophic forgetting.
3 Proposed Method
In this paper, we introduce a modular routing framework designed to address the stability-plasticity dilemma in MLLMs Shi et al. (2024). While modular architectures have demonstrated success in zero-shot generalization for text-only LLMs Muqeeth et al. (2024); Ostapenko et al. (2024), adapting this paradigm to multimodal continual learning requires addressing cross-modal interference. Our core insight is the decoupling of representation learning (skill acquisition) from task identification (routing). By isolating these processes, we ensure that new capabilities are integrated without eroding the model’s existing cross-modal alignment.
As illustrated in Figure 2, our framework operates in three distinct phases for each sequential task: (1) Expert Specialization via LoRA, (2) Router Vector Training, and (3) Dynamic Inference.
Figure 2: The proposed method. The process is divided into three steps: (1) Training task-specific LoRA adapters; (2) Training lightweight Routing Vectors (RV) to recognize task-specific activation patterns; (3) Inference using dynamic composition to select the top-$k$ experts per token.
3.1 Architecture Overview
We utilize a standard MLLM architecture composed of a vision encoder, a projection layer, and an LLM. To isolate the impact of our routing mechanism and ensure a controlled evaluation of catastrophic forgetting, we keep the vision encoder and projection layer frozen. We implement task-specific LoRA modules within the linear layers of the LLM.
3.2 Training Strategy
Our training strategy separates the acquisition of task knowledge from the task selection. This separation is critical for preventing the gradient conflict typically observed in sequential fine-tuning.
Step 1: Expert Specialization (LoRA).
@@ 3.2 Training Strategy
For each arriving task $T_{i}$, we specialize the LLM by training a dedicated LoRA adapter. Let $W_{0}\in\mathbb{R}^{d\times d_{in}}$ denote the frozen weights of a linear layer in the base LLM. For an input activation $u_{t}\in\mathbb{R}^{d_{in}}$ at token $t$, the output $h_{t}$ is computed as:
$h_{t}=W_{0}u_{t}+B_{i}A_{i}u_{t},$ (1)
where $A_{i}\in\mathbb{R}^{r\times d_{in}}$ and $B_{i}\in\mathbb{R}^{d\times r}$ are the trainable LoRA parameters with rank $r\ll d$. These parameters are optimized using only the dataset for task $T_{i}$, ensuring the expert captures task-specific nuances without being regularized by prior tasks.
Step 2: Routing Vector (RV) Training.
Once the expert for task $T_{i}$ is trained, we freeze parameters $A_{i}$ and $B_{i}$ and add a lightweight routing vector $v_{i}\in\mathbb{R}^{d_{in}}$. This vector serves as a soft gate, learning to identify the activation subspace associated with task $T_{i}$.
The routing vector is initialized to zero and optimized using the same next-token prediction objective as Step 1. During this phase, the layer output is governed by:
$h_{t}=W_{0}u_{t}+\sigma(v_{i}^{T}u_{t})\cdot(B_{i}A_{i}u_{t}),$ (2)
where $\sigma(\cdot)$ is the sigmoid function. By optimizing $v_{i}$ while the expert is active, the router learns a latent “signature” of the task within the activation space. Crucially, this step requires no access to data or modules from prior tasks.
3.3 Inference
Step 3: Dynamic Inference.
During inference, we utilize a bank of trained experts and their corresponding routing vectors $\mathcal{V}=\{v_{1},\dots,v_{N}\}$ for $N$ accumulated tasks. As depicted in Step 3 of Figure 2, routing decisions are computed dynamically at the token level, allowing the model to switch contexts mid-generation.
@@ 3.3 Inference
For an input activation $u_{t}$, we compute an affinity score with each routing vector:
$s_{t,j}=v_{j}^{T}u_{t},\quad j\in\{1,\dots,N\}.$ (3)
We then identify the set of indices $\mathcal{E}_{t}$ corresponding to the top-$k$ affinity scores. To effectively mix the selected experts, we normalize these scores using a softmax function with a scaling temperature $\tau$:
$w_{t,j}=\frac{\exp(s_{t,j}/\tau)}{\sum_{m\in\mathcal{E}_{t}}\exp(s_{t,m}/\tau)},\quad\forall j\in\mathcal{E}_{t}.$ (4)
The final output is the superposition of the base model and the weighted experts:
$h_{t}=W_{0}u_{t}+\sum_{j\in\mathcal{E}_{t}}w_{t,j}(B_{j}A_{j}u_{t}).$ (5)
This design ensures that the model dynamically retrieves the most relevant expertise for each token without requiring explicit task identifiers. This substantially mitigates catastrophic forgetting while allowing for positive transfer, as vision-centric tokens can route to vision experts and logic-centric tokens to reasoning experts.
4 Experiments
This section details our experimental setup, including model architectures, datasets, and training hyperparameters. We evaluate our routing framework on InternVL-2 Chen et al. (2024b) (2B and 8B variants) and LLaVA-v1.5-7B Liu et al. (2023) to assess scalability and robustness across diverse model families.
Table 1: Comparison of continual learning methods on InternVL2-2B and 8B. Bold indicates best performance; underline indicates second best. Performance is reported across four sequential specialization tasks and three general-purpose benchmarks.
Size Method COCO SNLI HM MMIMDb MMB ChartQA DocVQA
2B Zero-shot 0.795 52.9 60.2 46.6 79.1 59.5 84.7
Seq-FT 1.181 61.0 49.7 62.3 67.1 41.2 70.0
Arith. Merge 1.171 73.3 64.5 64.9 78.6 59.2 84.1
MR-LoRA 1.310 80.0 69.1 61.4 - - -